Parth (parthkohli):

\[1 + \frac{1}{4} + \frac{1\cdot 3}{4\cdot 8 } + \frac{1\cdot 3 \cdot 5}{4\cdot 8 \cdot 12} + \cdots\]

Parth (parthkohli):

@ganeshie8

OpenStudy (seascorpion1):

Your series simplifies to:\[\frac{(2n+1)!!}{(4n)!!}\] for n >=0

Parth (parthkohli):

what does it converge to

OpenStudy (seascorpion1):

I'm a bit rusty on this topic but I'll be back in 5mins with an explanation.

OpenStudy (dls):

1/2 ?

Parth (parthkohli):

nahi re, the first term itself is greater than 1/2.

OpenStudy (dls):

That is what @seascorpion1 's simplification converge to

OpenStudy (dls):

3/2 ?

Parth (parthkohli):

No, the answer is \(\sqrt 2\). You should recognise a Maclaurin series in there.

OpenStudy (seascorpion1):

Sorry I couldn't help in time but I learned from your answer.

OpenStudy (seascorpion1):

Are you sure, I get\[\frac{5\sqrt[8]{e}}{4}\]when I used wolfram to check it.

OpenStudy (dls):

I guess they are almost the same..accurate upto 2 decimal places

OpenStudy (dan815):